11. Problem 9-11 Problem Walk-Through Problem 9-11 Cost of Equity Radon Homes current EPS is $7.82. It was 4.66 5 years ago. The company pays out 40% of its earnings as dividends, and the stock sells for $36 a. Calculate the historical growth rate in earmings. (Hint: This is a 5-year growth period.) Do not round intermediate calculations. Round your answer to two decimal places. b. Calculate the next expected dividend per share, D1. (Hint: Do 0.40$7.82)-$3.13.) Assume that the past growth rate will continue. Do not round intermediate calculations. Round your answer to the nearest cent c. What is Radons cost of equity, r,? Do not round intermediate calculations. Round your answer to two decimal places. Continue without saving
Answer #1

A: Using financial calculator

Input: FV = 7.82

PV = 4.66

N= 5

Find I/Y as 10.91

Hence growth rate is 10.91%

B: D0= EPS*40% = 7.82*40%=3.13

D1 = D0*(1+g)

= 3.13*(1+0.1091)

=$3.47

C: Share price = D1/(Rs-g)

36= 3.47/ (Rs- 0.1091)

Rs= 20.55%